OpenStudy (anonymous):

Find the 23rd term of -12, -15, -18.

OpenStudy (jdoe0001):

do you know what the common ratios is?

OpenStudy (anonymous):

The difference is about -3?

OpenStudy (anonymous):

So it's an arithmetic series?

OpenStudy (jdoe0001):

ok so the common ratio is -3 the 1st term as you can see, is -12 so now that we know those two fellows we'll use them in \(\Large \bf a_{\color{red}{ n}}=a_1+({\color{red}{ n}}-1){\color{blue}{ d}} \\ \quad \\ a_1=\textit{first term}\\ {\color{blue}{ d}}=\textit{common ratio} \\ \quad \\ \Large a_{\color{red}{ 23}}=-12+({\color{red}{ 23}}-1){\color{blue}{ -3}}\)

OpenStudy (anonymous):

-12 + 22 - 3 = 7?

OpenStudy (jdoe0001):

well... the 22 * -3

OpenStudy (anonymous):

Oh I read that as subtract 3.

OpenStudy (anonymous):

-78?

OpenStudy (jdoe0001):

hmm my bad .... \(\large a_{\color{red}{ n}}=a_1+({\color{red}{ n}}-1){\color{blue}{ d}} \\ \quad \\ a_1=\textit{first term}\\ {\color{blue}{ d}}=\textit{common ratio} \\ \quad \\ \large a_{\color{red}{ 23}}=-12+({\color{red}{ 23}}-1)({\color{blue}{ -3}})\)

OpenStudy (jdoe0001):

-78? \(\Large \checkmark\)

OpenStudy (anonymous):

So for other problems, I just find the difference and plug in the values?

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

Thanks :)

OpenStudy (jdoe0001):

yw